sonu is 20 years older than monu. in 8 years sonu will be 4 times old as monu find thiwe present ages.

Answer by stanbon(75887)

sonu is 20 years older than monu. in 8 years sonu will be 4 times old as monu find thiwe present ages.
s = m + 20
(s+8) = 4(m+8)
----
substitute for "s" and solve for "m":
m + 20 + 8 = 4m + 32
3m = -4
-----
Note: 3m cannot be negative.
Check your problem statement.
Cheers,
Stan H.